# A Survey of World Models for Autonomous Driving

### v3

@@ 3.1 Generation of Future Physical World
GenAD[22] ECCV’24 Transformer Image + BEV Map + Trajectory nuScenes[36]
@@
CarFormer[72] ECCV’24 Transformer BEV + Trajectory + Destination + Traffic sign BEV + Action Longest6[73]
3.1.1 Image-based Generation
Dreamer Series. Dreamer series was first proposed for learning world models to address dynamic scene understanding in gaming and robotics [74, 75, 76], later expanded to autonomous driving with DriveDreamer [10]. It enables controllable video generation in high-density urban traffic using real-world driving data, though they remain limited to 2D outputs lacking spatial-temporal coherence. Subsequent research introduces DriveDreamer-2 [37], incorporating LLM-based prompting mechanisms to enhance interactivity and diversity. With further advancements, DriveDreamer4D [26] significantly extends model capabilities by utilizing world model priors to generate spatial-temporally coherent 4D driving videos, marking a transition from traditional 2D to advanced 4D video generation. ReconDreamer [39] explores online restoration techniques to achieve precise video reconstruction of dynamic scenes. WorldDreamer [40] expands the applicability of world models by pioneering masked token prediction for multi-modal generation, bridging text-to-video and action-to-video tasks, and moving towards general-purpose world model construction. The development of the Dreamer series demonstrates a clear trend: progressing from 2D to multi-dimensional video generation, transitioning from task-specific to general-purpose models, and shifting from closed, structured conditions toward more open and general natural-language-based conditions.
Diffusion-based Image Generation. The evolution of this type in autonomous driving are coalescing around two complementary thrusts: controllable multi-view generation (e.g., BEVControl [43] and DrivingDiffusion [14]), which use latent diffusion to render geometry-consistent videos from rich cues such as BEV layouts, text prompts and optical flow, and high-fidelity spatio-temporal modelling, where Drive-WM [44] and Vista [46] push diffusion roll-outs to longer horizons and higher spatial resolutions, markedly boosting temporal coherence and per-frame detail. The field progresses through breakthroughs in multi-modal conditions: integrating ego trajectory, human-pose, and DINO appearance tokens for behaviour-aware video synthesis [16]; translating simulator layouts into photorealistic scenes through a cascade diffusion pipeline [47] or dual-scale controlNet-stable diffusion [77]; uniting LiDAR and camera streams in a BEV latent grid for geometry-consistent cross-sensor forecasting [32]; aligning ego–other vehicle trajectories in latent space to enable fully controllable, multi-agent video generation [54]; and grounding multi-camera driving video generation in physics by aligning ego-world coordinates, injecting 3D flows, and box-guided occlusion reasoning [50]. Recent methods [55, 45, 56, 57, 58, 59, 48] focus on expanding from simple RGB image generation to more diverse and enriched paradigms, including extending conventional RGB prediction to jointly generate both images and depth [55]; shifting from layout-conditioned synthesis to fine-grained trajectory-conditioned control [45]; moving beyond 2D-only conditioning toward multi-view consistent generation [56]; transitioning from 2D layout cues to full 3D scene conditioned synthesis [57]; broadening single-view synthesis into multi-view, multi-agent interactive scenarios [58]; progressing from short-term clips to long-horizon video generation [59]; and further leveraging failure-case augmentation for robustness [48]. This technical evolution reveals a clear trend from isolated scene generation to integrated closed-loop simulation systems, providing high-fidelity, scalable synthetic environments for autonomous driving.
Transformer-based Image Generation. Transformer-based driving imagers are evolving from narrow, single-view renderers into holistic, action-aware world models. Cross-view architectures such as HoloDrive [61] and BEVGen [62] fuse camera and LiDAR or lift BEV layouts to street-level frames, unifying 2D appearance with 3D geometry. A parallel line scales sequence modeling: DrivingWorld [64] and GAIA-1 [15] treat video, text and control as one token stream, enabling minute-long, controllable clips with coherent semantics. These token factories prove that unsupervised prediction alone can encode road rules and actor intent. Complementary work probes limits and remedies: DriveSim [66] exposes causal hallucinations in multimodal LLMs, while a DINOv2-based multi-view encoder [35] compresses synchronized front views into a latent scene grid that decodes to both RGB and BEV semantics, reducing covariate shift.
@@ 3.1.1 Image-based Generation
Summary. Table I shows different modern image-centric generative frameworks, which use diffusion and transformer architectures to produce diverse driving data. Several observations can be drawn:
[…]
@@
3.1.2 BEV-based Generation
BEV Representation. A BEV representation unifies multi-modal sensor streams into a single top-down map, exposing lane layouts, dynamic actors, and free space in a planner-friendly format [32]. By stripping away perspective distortions, BEV methods provide high-level, spatially coherent context that has become a powerful complement – or even alternative – to image-centric pipelines, supporting tasks such as motion prediction and long-horizon trajectory forecasting. This 2D projection, however, gains simplicity at the expense of depth resolution: BEV maps can struggle to retain fine-grained 3D geometry in scenes with complex vertical structure or steep depth gradients [78].
Starting with the probabilistic BEV forecaster FIERY [33], researchers stretch temporal range with StretchBEV [70], learn continuous 4D occupancy fields with UnO [31], and cut computation via the 2DCNN design [71]. Object-centric slots then emerge: CarFormer [72] assigns every actor a latent slot. MILE [34] compresses the whole scene into a single compact latent with the driving policy, enabling policy-aware, multi-modal prediction. Together, these advances unlock generative planning; GenAD [22] rolls the latent forward to sample ego-conditioned futures, BEVControl [43] lets users overwrite it with editable sketches for safety auditing, and ViDAR [20], although operating in point-cloud space, supplies geometry-aware pre-training that lifts BEV performance across the board. Together these works chart a path from early probabilistic forecasts to efficient, object-aware, user-controllable world models.
Summary. Table I summarizes recent BEV‑based generative methods, which predict future BEV representations for autonomous driving tasks. Several observations can be drawn:
[…]
@@ 3.1.2 BEV-based Generation
TABLE II: Overview of summary, methods and core architectures for OG-/PC-based generation (§3.1.3 and §3.1.4).
@@
OG-based Generation CNNs Occ4cast[79] IROS’24 CNNs LiDAR Occupancy OCFBench[79]
Cam4DOcc[80] CVPR’24 CNN Image + Ego motion + GMO + GSO Occupancy OpenOccupancy[81]
PreWorld[82] ICLR’25 CNN Image + Occupancy Occupancy Occ3D[83]
[…]
OccWorld[85] ECCV’24 Transformer Occupancy + Trajectory Occupancy + Trajectory Occ3D[83]
DFIT-OccWorld[86] arXiv’24 Transformer Occupancy + Image + Ego poses Occupancy + Image + Trajectory Occ3D[83] + OpenScene[87]
DriveWorld[21] CVPR’24 Transformer Image + Action Occupancy + Action nuScenes[36] + OpenScene[87]
Drive-OccWorld[88] AAAI’25 Transformer Image + Action Occupancy + Flow + Trajectory nuScenes[36] + OpenOccupancy[81]
OccLLaMA[89] arXiv’24 Transformer Occupancy + Trajectory + Text Occupancy + Trajectory + Text nuScenes[36] + OpenOccupancy[81]
GaussianWorld[90] CVPR’25 Transformer + 3DGS Gaussians + Image Occupancy nuScenes[36]
RenderWorld[91] ICRA’24 Transformer + 3DGS Image Occupancy + Trajectory nuScenes[36]
OccProphet[92] ICLR’25 Transformer Image Occupancy + Flow nuScenes[36] + OpenOccupancy[81]
Occ-LLM[93] ICRA’25 LLM Image + Occupancy Occupancy + Trajectory + Text nuScenes[36] + Occ3D[83]
T3Former [94] arXiv’25 Transformer Occupancy Occupancy + Trajectory nuScenes[36]
[…]
OccSora[18] arXiv’24 Diffusion Model Occupancy + Trajectory Occupancy Occ3D[83]
UniScene[96] arXiv’24 Diffusion Model BEV Layouts + Text Occupancy + LiDAR + Image OpenOccupancy[81] + nuScenes[36]
DynamicCity[17] ICLR’25 Diffusion Model Occupancy + Command + Trajectory + Layout Occupancy Occ3D[83] + CarlaSC[97]
@@
PC-based Generation CNNs PCP[98] CoRL’22 CNN LiDAR + Range map LiDAR KITTI[99]
@@
4DOcc[100] CVPR’23 CNN LiDAR Occupancy + LiDAR KITTI[99] + nuScenes[36]
[…]
@@
PC-based Generation Transformers ViDAR[20] CVPR’24 Transformer Image + Action LiDAR nuScenes[36]
@@
HERMES[102] arXiv’25 Transformer Image + Text + Action LiDAR + Text nuScenes[36]
[…]
@@
Copilot4D[105] ICLR’24 Diffusion Model LiDAR LiDAR KITTI[99] + nuScenes[36]
@@
RangeLDM[106] arXiv’24 Diffusion Model LiDAR + Range map LiDAR KITTI-360[104]
[…]
lidarGeneration[107] IROS’19 VAE + GAN LiDAR + Point map LiDAR KITTI[99]
@@
UltraLiDAR[108] CVPR’23 VQ-VAE LiDAR LiDAR KITTI-360[104] + nuScenes[36]
@@
SPFNet[109] CoRL’21 LSTMs LiDAR + Range map LiDAR KITTI[99] + nuScenes[36]
@@
S2net[110] ECCV’22 LSTMs LiDAR + Range map LiDAR KITTI[99] + nuScenes[36]
NFL[111] ICCV’23 Nerf LiDAR LiDAR KITTI[99]
@@
Nerf-lidar[112] AAAI’24 Nerf LiDAR + Image + Segmentation LiDAR nuScenes[36]
3.1.3 OG-based Generation
OG Representation. An OG representation divides the driving scene into 3D voxels and assigns each cell a probability of being occupied, producing a single lattice that simultaneously tracks static structure and moving actors. This discretisation yields far richer geometric detail than 2D projections such as BEV, enabling fine-grained reasoning about scene evolution and actor interactions – qualities that have made OGs a mainstay of future-state prediction in autonomous driving. The price of this fidelity, however, is steep: dense voxel grids demand considerable memory and computational throughput, which can hinder deployment in large-scale or real-time systems [83].
OG-based Generation (i.e., occupancy forecasting) originates from forecasting semantic occupancy grids on BEV [113, 114, 28, 115, 116, 117]. It is employed to predict how the surrounding occupancy will evolve in the near future.
CNN-based OG Generation. Occ4cast starts this type by unifying sparse-to-dense completion and 4D forecasting into a coherent Eulerian grid that reconstructs geometry and predicts its future [79]. Cam4DOcc then ports this paradigm from LiDAR to low-cost multi-view cameras, releasing a benchmark and network that align per-view geometry with motion cues [80]. Finally, PreWorld trims supervision demands by introducing a semi-supervised, two-stage pipeline that distills 2D imagery into 3D occupancy and planning priors [82]. These milestones trace a clear trajectory: from dense LiDAR-based geometry, through camera-centric low-cost sensing, to label-efficient world models.
Transformer-based OG Generation. ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ archi- ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ open‑loop ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ “Interaction between Planning and Prediction”.
@@ 3.1.3 OG-based Generation
Transformer-based OG research starts with MUVO [84], which fuses camera + LiDAR features in a voxel-level Transformer to predict 3D occupancy, yet its deterministic rollout and label hunger limit scalability. OccWorld [85] converts these voxels into autoregressive scene tokens, while DFIT-OccWorld [86] decouples dynamic-static warping to cut training cost, collectively launching a generative, token-based line of work. Toward perception-control unity, DriveWorld [21] pre-trains a memory-augmented world model for downstream planning, and Drive-OccWorld [88] adds trajectory-conditioned heads so the same tokens directly score driving costs. Efficiency remains a parallel thread – OccProphet [92] refines compute, and T3Former [94] sparsifies temporal attention with triplane features, achieving real-time, camera-only deployment. OccLLaMA [89] and Occ-LLM [93] broaden the token vocabulary to language and action, enabling instruction-driven occupancy reasoning across the same generative backbone. Collectively, the field has evolved from deterministic voxel Transformers through autoregressive token worlds and efficiency-focused triplane models to multimodal, instruction-aware, real-time world models. Some studies attempt to introduce 3D Gaussian Splatting (3DGS) into this field. GaussianWorld [90] treats occupancy as a 4D forecasting task in Gaussian space, explicitly factorising ego motion, local dynamics and scene completion; RenderWorld [91] removes all lidar by pairing a self‑supervised Img2Occ labeler with an AM‑VAE encoder and the same autoregressive backbone, yielding a compact, vision‑only pipeline. Overall, these studies trace a path from heavy deterministic voxel fusion to lightweight stochastic tokenised representations, from perception‑only forecasting to memory‑ and planning‑conditioned roll‑outs, and from single‑modality grids to language‑grounded multi‑sensor inputs; the overall trend converges on controllable, efficient, multi‑modal world models.
${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ gene-
@@
${}_{\!}$ methods enable fine‑grained control by coupling spatio‑temporal diffusion transformers with compact 4D tokens. They further advance toward controllable, token-efficient, multi-modal world models by: introducing trajectory-resampling control to steer future occupancy synthesis [95]; compressing scenes into compact 4D tokens for coherent occupancy videos [18]; projecting 4D voxels onto HexPlanes and rolling them out with DiT for large-scale semantic forecasting [17]; and diffusing unified BEV semantic grids into joint video + LiDAR streams [96].
@@
Summary. Table II assembles recent OG generative frameworks using CNNs, Transformers, and diffusion models for 4D occupancy forecasting. Several observations emerge:
[…]
    ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ cost.
[…]

### v4

@@ 3.1.1. Image-based Generation
GenAD(Zheng et al., 2025b) ECCV’24 Transformer Image + BEV Map + Trajectory nuScenes(Caesar et al., 2020)
@@
CarFormer(Hamdan and Güney, 2025) ECCV’24 Transformer BEV + Trajectory + Destination + Traffic sign BEV + Action Longest6(Chitta et al., 2022)
@@
Summary. Table 1 shows different modern image-centric generative frameworks, which use diffusion and transformer architectures to produce diverse driving data. Several observations can be drawn:
@@
3.1.2. BEV-based Generation
BEV Representation. A BEV representation unifies multi-modal sensor streams into a single top-down map, exposing lane layouts, dynamic actors, and free space in a planner-friendly format (Zhang et al., 2024a). By stripping away perspective distortions, BEV methods provide high-level, spatially coherent context that has become a powerful complement – or even alternative – to image-centric pipelines, supporting tasks such as motion prediction and long-horizon trajectory forecasting. This 2D projection, however, gains simplicity at the expense of depth resolution: BEV maps can struggle to retain fine-grained 3D geometry in scenes with complex vertical structure or steep depth gradients (Feng et al., 2023b).
Starting with the probabilistic BEV forecaster FIERY (Hu et al., 2021b), researchers stretch temporal range with StretchBEV (Akan and Güney, 2022), learn continuous 4D occupancy fields with UnO (Agro et al., 2024), and cut computation via the 2DCNN design (Li et al., 2023). Object-centric slots then emerge: CarFormer (Hamdan and Güney, 2025) assigns every actor a latent slot. MILE (Hu et al., 2022b) compresses the whole scene into a single compact latent with the driving policy, enabling policy-aware, multi-modal prediction. Together, these advances unlock generative planning; GenAD (Zheng et al., 2025b) rolls the latent forward to sample ego-conditioned futures, BEVControl (Yang et al., 2023b) lets users overwrite it with editable sketches for safety auditing, and ViDAR (Yang et al., 2024a), although operating in point-cloud space, supplies geometry-aware pre-training that lifts BEV performance across the board. Together these works chart a path from early probabilistic forecasts to efficient, object-aware, user-controllable world models.
Summary. Table 1 summarizes recent BEV‑based generative methods, which predict future BEV representations for autonomous driving tasks. Several observations can be drawn:
@@ 3.1.2. BEV-based Generation
3.1.3. OG-based Generation
OG Representation. An OG representation divides the driving scene into 3D voxels and assigns each cell a probability of being occupied, producing a single lattice that simultaneously tracks static structure and moving actors. This discretisation yields far richer geometric detail than 2D projections such as BEV, enabling fine-grained reasoning about scene evolution and actor interactions – qualities that have made OGs a mainstay of future-state prediction in autonomous driving. The price of this fidelity, however, is steep: dense voxel grids demand considerable memory and computational throughput, which can hinder deployment in large-scale or real-time systems (Tian et al., 2024).
OG-based Generation (i.e., occupancy forecasting) originates from forecasting semantic occupancy grids on BEV (Mahjourian et al., 2022; Schreiber et al., 2019; Casas et al., 2021; Sadat et al., 2020; Toyungyernsub et al., 2022; Khurana et al., 2022). It is employed to predict how the surrounding occupancy will evolve in the near future.
CNN-based OG Generation. Occ4cast starts this type by unifying sparse-to-dense completion and 4D forecasting into a coherent Eulerian grid that reconstructs geometry and predicts its future (Liu et al., 2024). Cam4DOcc then ports this paradigm from LiDAR to low-cost multi-view cameras, releasing a benchmark and network that align per-view geometry with motion cues (Ma et al., 2024a). Finally, PreWorld trims supervision demands by introducing a semi-supervised, two-stage pipeline that distills 2D imagery into 3D occupancy and planning priors (Li et al., 2025b). These milestones trace a clear trajectory: from dense LiDAR-based geometry, through camera-centric low-cost sensing, to label-efficient world models.
Table 2. Overview of summary, methods and core architectures for OG-/PC-based generation (§3.1.3 and §3.1.4).
@@ 3.1.3. OG-based Generation
OG-based Generation CNNs Occ4cast(Liu et al., 2024) IROS’24 CNNs LiDAR Occupancy OCFBench(Liu et al., 2024)
Cam4DOcc(Ma et al., 2024a) CVPR’24 CNN Image + Ego motion + GMO + GSO Occupancy OpenOccupancy(Wang et al., 2023)
PreWorld(Li et al., 2025b) ICLR’25 CNN Image + Occupancy Occupancy Occ3D(Tian et al., 2024)
[1pt/1pt] OG-based Generation Transformers MUVO(Bogdoll et al., 2023) IV’23 Transformer Image + LiDAR + Action Image + LiDAR + Occupancy CARLA(Dosovitskiy et al., 2017a)
OccWorld(Zheng et al., 2025a) ECCV’24 Transformer Occupancy + Trajectory Occupancy + Trajectory Occ3D(Tian et al., 2024)
DFIT-OccWorld(Zhang et al., 2024c) arXiv’24 Transformer Occupancy + Image + Ego poses Occupancy + Image + Trajectory Occ3D(Tian et al., 2024) + OpenScene(Contributors, 2023)
DriveWorld(Min et al., 2024b) CVPR’24 Transformer Image + Action Occupancy + Action nuScenes(Caesar et al., 2020) + OpenScene(Contributors, 2023)
Drive-OccWorld(Yang et al., 2025c) AAAI’25 Transformer Image + Action Occupancy + Flow + Trajectory nuScenes(Caesar et al., 2020) + OpenOccupancy(Wang et al., 2023)
OccLLaMA(Wei et al., 2024) arXiv’24 Transformer Occupancy + Trajectory + Text Occupancy + Trajectory + Text nuScenes(Caesar et al., 2020) + OpenOccupancy(Wang et al., 2023)
GaussianWorld(Zuo et al., 2025) CVPR’25 Transformer + 3DGS Gaussians + Image Occupancy nuScenes(Caesar et al., 2020)
RenderWorld(Yan et al., 2025a) ICRA’24 Transformer + 3DGS Image Occupancy + Trajectory nuScenes(Caesar et al., 2020)
OccProphet(Chen et al., 2025b) ICLR’25 Transformer Image Occupancy + Flow nuScenes(Caesar et al., 2020) + OpenOccupancy(Wang et al., 2023)
Occ-LLM(Xu et al., 2025a) ICRA’25 LLM Image + Occupancy Occupancy + Trajectory + Text nuScenes(Caesar et al., 2020) + Occ3D(Tian et al., 2024)
T3Former (Xu et al., 2025b) arXiv’25 Transformer Occupancy Occupancy + Trajectory nuScenes(Caesar et al., 2020)
AFMWM(Liu et al., 2025) arXiv’25 Transformer Occupancy + Trajectory Occupancy + Video Occ3D(Tian et al., 2024) + KITTI360(Liao et al., 2022)
I2-World (Liao et al., 2025) arXiv’25 Transformer Occupancy 4D occupancy Occ3D(Tian et al., 2024)
[1pt/1pt] OG-based Generation Diffusion Models DOME(Gu et al., 2024) arXiv’24 Diffusion Model Occupancy + Action Occupancy nuScenes(Caesar et al., 2020)
OccSora(Wang et al., 2024b) arXiv’24 Diffusion Model Occupancy + Trajectory Occupancy Occ3D(Tian et al., 2024)
UniScene(Li et al., 2024a) arXiv’24 Diffusion Model BEV Layouts + Text Occupancy + LiDAR + Image OpenOccupancy(Wang et al., 2023) + nuScenes(Caesar et al., 2020)
DynamicCity(Bian et al., 2025) ICLR’25 Diffusion Model Occupancy + Command + Trajectory + Layout Occupancy Occ3D(Tian et al., 2024) + CarlaSC(Wilson et al., 2022)
COME(Shi et al., 2025b) arXiv’25 Diffusion Model Occupancy + Pose/BEV layout Occupancy Occ3D(Tian et al., 2024)
@@
PC-based Generation CNNs PCP(Mersch et al., 2022) CoRL’22 CNN LiDAR + Range map LiDAR KITTI(Geiger et al., 2012)
@@
4DOcc(Khurana et al., 2023) CVPR’23 CNN LiDAR Occupancy + LiDAR KITTI(Geiger et al., 2012) + nuScenes(Caesar et al., 2020)
[1pt/1pt] PCPNet(Luo et al., 2023) RA-L’23 CNN LiDAR + Range map LiDAR KITTI(Geiger et al., 2012)
@@
PC-based Generation Transformers ViDAR(Yang et al., 2024a) CVPR’24 Transformer Image + Action LiDAR nuScenes(Caesar et al., 2020)
@@
HERMES(Zhou et al., 2025) ICCV’25 Transformer Image + Text + Action LiDAR + Text nuScenes(Caesar et al., 2020)
[1pt/1pt] PC-based Generation Diffusion Models LiDARGen(Zyrianov et al., 2022) ECCV’22 Diffusion Model LiDAR + Range map LiDAR KITTI-360(Liao et al., 2022)
@@
Copilot4D(Zhang et al., 2024b) ICLR’24 Diffusion Model LiDAR LiDAR KITTI(Geiger et al., 2012) + nuScenes(Caesar et al., 2020)
@@
RangeLDM(Hu et al., 2024b) arXiv’24 Diffusion Model LiDAR + Range map LiDAR KITTI-360(Liao et al., 2022)
LiDARCrafter(Liang et al., 2025a) arXiv’25 Diffusion Model Text + Ego-centric scene graph + LiDAR + Motion LiDAR nuScenes(Caesar et al., 2020)
[1pt/1pt] PC-based Generation Others Lidarsim(Manivasagam et al., 2020) CVPR’20 Composition LiDAR LiDAR Internal dataset(Manivasagam et al., 2020)
lidarGeneration(Caccia et al., 2019) IROS’19 VAE + GAN LiDAR + Point map LiDAR KITTI(Geiger et al., 2012)
@@
UltraLiDAR(Xiong et al., 2023) CVPR’23 VQ-VAE LiDAR LiDAR KITTI-360(Liao et al., 2022) + nuScenes(Caesar et al., 2020)
@@
SPFNet(Weng et al., 2021) CoRL’21 LSTMs LiDAR + Range map LiDAR KITTI(Geiger et al., 2012) + nuScenes(Caesar et al., 2020)
@@
S2net(Weng et al., 2022) ECCV’22 LSTMs LiDAR + Range map LiDAR KITTI(Geiger et al., 2012) + nuScenes(Caesar et al., 2020)
NFL(Huang et al., 2023a) ICCV’23 Nerf LiDAR LiDAR KITTI(Geiger et al., 2012)
@@
Nerf-lidar(Zhang et al., 2024d) AAAI’24 Nerf LiDAR + Image + Segmentation LiDAR nuScenes(Caesar et al., 2020)
Transformer-based OG Generation. Many recent leading methods of this type adopt Transformer‑based core architectures. These models act as vision‑centric world models for autonomous driving and are designed to learn from pre‑collected, open‑loop benchmarks; hence they sit outside “Interaction between Planning and Prediction”.
@@
Transformer-based OG research starts with MUVO (Bogdoll et al., 2023), which fuses camera + LiDAR features in a voxel-level Transformer to predict 3D occupancy, yet its deterministic rollout and label hunger limit scalability. OccWorld (Zheng et al., 2025a) converts these voxels into autoregressive scene tokens, while DFIT-OccWorld (Zhang et al., 2024c) decouples dynamic-static warping to cut training cost, collectively launching a generative, token-based line of work. Toward perception-control unity, DriveWorld (Min et al., 2024b) pre-trains a memory-augmented world model for downstream planning, and Drive-OccWorld (Yang et al., 2025c) adds trajectory-conditioned heads so the same tokens directly score driving costs. Efficiency remains a parallel thread – OccProphet (Chen et al., 2025b) refines compute, and T3Former (Xu et al., 2025b) sparsifies temporal attention with triplane features, achieving real-time, camera-only deployment. Meanwhile, AFMWM(Liu et al., 2025) pairs a high-ratio Swin-VAE compressor with efficient latent forecasting. I2-World (Liao et al., 2025) further pursues efficiency by decoupling tokenization into intra- and inter-scene stages within an encoder–decoder with a transformation-matrix control head. OccLLaMA (Wei et al., 2024) and Occ-LLM (Xu et al., 2025a) broaden the token vocabulary to language and action, enabling instruction-driven occupancy reasoning across the same generative backbone. Collectively, the field has evolved from deterministic voxel Transformers through autoregressive token worlds and efficiency-focused triplane models to multimodal, instruction-aware, real-time world models. Some studies attempt to introduce 3D Gaussian Splatting (3DGS) into this field. GaussianWorld (Zuo et al., 2025) treats occupancy as a 4D forecasting task in Gaussian space, explicitly factorising ego motion, local dynamics and scene completion; RenderWorld (Yan et al., 2025a) removes all lidar by pairing a self‑supervised Img2Occ labeler with an AM‑VAE encoder and the same autoregressive backbone, yielding a compact, vision‑only pipeline. Overall, these studies trace a path from heavy deterministic voxel fusion to lightweight stochastic tokenised representations, from perception‑only forecasting to memory‑ and planning‑conditioned roll‑outs, and from single‑modality grids to language‑grounded multi‑sensor inputs; the overall trend converges on controllable, efficient, multi‑modal world models.
@@
Diffusion-based OG Generation. Diffusion‑based OG generation methods enable fine‑grained control by coupling spatio‑temporal diffusion transformers with compact 4D tokens. They further advance toward controllable, token-efficient, multi-modal world models by: introducing trajectory-resampling control to steer future occupancy synthesis (Gu et al., 2024); compressing scenes into compact 4D tokens for coherent occupancy videos (Wang et al., 2024b); projecting 4D voxels onto HexPlanes and rolling them out with DiT for large-scale semantic forecasting (Bian et al., 2025); diffusing unified BEV semantic grids into joint video + LiDAR streams (Li et al., 2024a); and disentangling ego-motion from scene evolution via a scene-centric forecasting branch whose guidance is injected through a tailored ControlNet into the occupancy world model (Shi et al., 2025b).
@@
Summary. Table 2 assembles recent OG generative frameworks using CNNs, Transformers, and diffusion models for 4D occupancy forecasting. Several observations emerge:
  • The field migrates from heavy, deterministic voxel fusion to lightweight, stochastic tokenisation and further to Gaussian splats, reducing memory cost.